OpenStudy (bluemoon1):

Y=-3/2x+1 find the slope of a line perpendicular to each given line

OpenStudy (steve816):

Remember that a perpendicular slope is the opposite reciprocal.

OpenStudy (bluemoon1):

I have no idea how to do that

OpenStudy (bluemoon1):

Like it would help if you gave me an equation

OpenStudy (amorfide):

if you have m=a then the perpendicular line will have a gradient of -1/m=-1/a which is known as the negative reciprocal

OpenStudy (amorfide):

for example, y=3x+2 the perpendicular line would have a slope (gradient) of -1/3

OpenStudy (steve816):

@amorfide Not exactly negative. If you have a negative number, than it turns into positive.

OpenStudy (bluemoon1):

So -1/3 would be the answer? It would be the slope?

OpenStudy (amorfide):

I never meant the answer will be negative, but it will be the negative reciprocal

OpenStudy (bluemoon1):

I thought you divide by 1 or something. So how does this turn into an equation that I can solve?

OpenStudy (amorfide):

in your question y=-3/2x +1 m=-3/2 so -1/m=-1/(-3/2)=-2/-3=2/3

OpenStudy (bluemoon1):

Do I solve that?

OpenStudy (amorfide):

you are only asked for the slope of the perpendicular line, so that is your answer

OpenStudy (bluemoon1):

Cool thanks!

OpenStudy (bluemoon1):

You just saved me so much time